OpenStudy (anonymous):

PLEASE HELP !! Choose the slope-intercept: equation of the line that passes through the point (6, 4) and is parallel to y = 1/3x - 3. A) y = 3x - 14 B) y = 1/3x + 2 C) y = -3x + 22 D) y = -1/3 x + 8

OpenStudy (zzr0ck3r):

y = -3x+b so you know its c

OpenStudy (zzr0ck3r):

4 = -3(6) + b b = 22 y = -3x+22

OpenStudy (zzr0ck3r):

any time we are looking for an orthogonal linear equation to some other equation take the negative of the reciprocal of the slop of the line we are going orthogonal to.

OpenStudy (zzr0ck3r):

(1/3) becomes -3

OpenStudy (anonymous):

oh... okay... would the same principles apply for this problem? Choose the slope-intercept: equation of the line that passes through the point (6, -6) and is perpendicular to y = 3x - 6.

OpenStudy (zzr0ck3r):

yeah

OpenStudy (zzr0ck3r):

we know it will be y = (-1/3)x+b then use the point -6= (-1/3)(6)+b -6 = -2+b b = -4 so y= (-1/3)x-4

OpenStudy (anonymous):

ohhhh

OpenStudy (zzr0ck3r):

the only information you need is the slop of the original line

OpenStudy (zzr0ck3r):

slope*